### v1

@@ 3 Numerical bifurcation analysis
To investigate the dynamical behaviour of the system beyond the equilibria, in this section we perform a numerical bifurcation analysis of the system for the parameter set considered by [20] and [6], which is plausible with a highly infectious disease like pertussis. We take $\mathcal{R}_{0}=15$, $\gamma=17$ (year-1), corresponding to an average infectious period of 21 days, and per capita death rate $d=0.02$ (year-1), corresponding to an average lifetime of 50 years. The parameter $\beta$ is then fixed through (2) ($\beta=255.3$). We perform stability and bifurcation analyses varying the parameters $\tau$ and $\nu$, which represent the maximal duration of immunity in the absence of boosting and the boosting force, respectively. Since $\mathcal{R}_{0}>1$, the endemic equilibrium exists for all $\nu$ and $\tau$, with the susceptible population equal to $S^{*}=\mathcal{R}_{0}^{-1}=\frac{1}{15}$ at equilibrium.
@@
Figure 2: Left: stability of the endemic equilibrium of model (1) in the plane $(\nu,\tau)$. Solid lines represent Hopf bifurcation curves, with HH ($\bullet$) and GH ($\circ$) bifurcations detected along the curves. The endemic equilibrium is unstable in the shaded area. Although not visible in the diagram due to the scaling of the axis, the lowest curves connect at a large value of $\nu$. The vertical dashed lines correspond to the parameters chosen for the one-parameter bifurcation diagrams in the following analyses. Note that an accurate approximation near $\nu=0$ requires large discretization indices and becomes computationally challenging, so we do not show it here. Right: close-up with the bifurcations numerically detected on different branches of limit cycles, delimiting their stability regions.
@@
Figure 2 shows the computed stability regions of the endemic equilibrium in the parameter plane $(\nu,\tau)$, which matches with the results presented by Barbarossa et al. [6, Figure 1]. The endemic equilibrium is stable outside the outermost boundary curve, and unstable inside (shaded region). Note that, as theoretically expected, the equilibrium is always stable on the axis $\tau=0$ (corresponding to an SIS-type model with individuals immediately susceptible upon recovery).
@@
GH points are often associated with the emergence of limit points of cycles (LPC) [19, Section 8.3.3] and hence with bistability of equilibria and periodic solutions. Similarly, HH points are associated with the emergence of Neimark–Sacker bifurcations (NS) and invariant tori [19, Section 8.6.3]. Figure 2 therefore suggests the emergence of rich and complex dynamics. We further explore such complex dynamics by studying the one-parameter continuation of the periodic orbits emerging from Hopf and their bifurcations, fixing $\nu=4.8$, $3.2$, $2$, $1$, and varying the duration of immunity $\tau$ (along the dashed lines indicated in Figure 2).
Figure 3: Left: bifurcation diagram of $I$ varying $\tau$, for $\nu=4.8$, including the equilibrium curve with the detected Hopf bifurcations ($\bullet$) and the max/min values of the branch of limit cycles. Solid red lines indicate a stable equilibrium/limit cycle, dashed lines are unstable. Right: period of the limit cycle branch (years).
@@
Figure 3 shows the one-parameter bifurcation diagram with respect to $\tau$ for $\nu=4.8$ (left), and the corresponding estimated period along the branch of limit cycles. Four Hopf bifurcations are detected along the equilibrium curve, at approximately $\tau\approx 0.05$, $2.86$, $3.15$, and $4.59$. The equilibrium is unstable between the first two points and between the last two points. A bubble of stable limit cycles exists for $\tau$ between $3.15$ and $4.59$ (both supercritical bifurcations), with a period between 2.2 and 2.8 years. When computing the limit cycle branch emerging from $\tau\approx 0.05$, the amplitude of the branch diverges quickly, and the numerical continuation stops. This branch is therefore not shown in the figure.
@@
As we can observe in Figure 5, the numerical analysis, i.e., the rightmost eigenvalue crossings as $\tau$ varies (bottom left figure), aligns perfectly with the computation of possible stability switches, which are the zeros of $S^{\pm}_{n}(\tau)$, $n=0,1$ (top left figure). The zeros were calculated using (19) on the corresponding intervals $J^{+}$ and $J^{-}$. This also confirms Lemma 2, which implies that the endemic equilibrium will be stable for all $\tau>\tau_{max}$.
Figure 5: Top: graph of the stability switch functions $S_{n}^{\pm}(\tau)$ for $n=0,1,2$, where $\nu=4.8$ and $3.2$, respectively. Bottom: corresponding real part of the rightmost eigenvalues along the equilibrium branch, varying $\tau$. Black dots indicate Hopf bifurcations.
To illustrate the emergence of bistability through the GH point, Figure 6 shows the one-parameter continuation for $\nu=3.2$. In this case, we still observe a bubble of limit cycles for $\tau$ between the Hopf points $2.11$ and $5.37$ (first supercritical, second subcritical). An LPC and an NS bifurcations are detected on the branch of limit cycles, and the limit cycles are stable between those bifurcations. Bistability occurs for values of $\tau$ between the Hopf and LPC bifurcations (i.e., between $5.37$ and $6.16$), where a stable limit cycle coexists with the stable endemic equilibrium, so the asymptotic behaviour of the system depends on the initial condition.
@@
Figure 6: Left: bifurcation diagram of $I$ varying $\tau$, for $\nu=3.2$. Solid red lines indicate stable equilibrium/limit cycles; dashed lines indicate unstable. NS and LPC labels indicate Neimark-Sacker and limit point of cycles bifurcations, respectively. The blue shading shows the interval in which a stable equilibrium coexists with a stable limit cycle, for $\tau$ between $5.37$ and $6.16$. Right: period along the limit cycle branch (years).
@@
Figure 7: Left: bistability of the positive equilibrium and the stable limit cycle for $\nu=3.2$ and $\tau=5.6$. Right: emergence of a stable torus via an NS bifurcation, for $\nu=3.2$ and $\tau=4.2$.
An even more complex dynamical picture appears for $\nu=2$, see Figure 8, with two different closed limit cycle curves connecting different pairs of Hopf bifurcations, one with larger amplitude and larger period than the other. Each branch is stable in an interval of values between a LPC and a NS bifurcation. Two regions of bistability appear: one with a stable equilibrium coexisting with a stable limit cycle, and one with two different stable limit cycles. These regions are shaded in Figure 8. Figure 9 shows the orbits of the limit cycles in the smaller- and larger-amplitude branches, as well as their time profiles. The figure highlights the differences in both amplitude and period. Following the LPC and NS curves for each bubble of limit cycles in the plane $(\nu,\tau)$ one can get some insights about regions of parameters where two stable periodic solutions coexist, or one stable periodic solution coexists with the stable endemic equilibrium. These curves are sketched in the right panel in Figure 2.
Figure 8: Left: bifurcation diagram of $I$ varying $\tau$, for $\nu=2$. Solid red lines indicate stable equilibrium/limit cycles; dashed indicate unstable. Right: period of the limit cycle along the branches. The figure shows two intervals of bistability: the blue shading indicates bistability of an equilibrium and limit cycle, while the red shading indicates bistability of two limit cycles with different amplitude and period.
Figure 9: Orbits (left) and trajectories (right) of the stable coexisting limit cycles, for $\nu=2$ and $\tau=7$. Periods are approximately $1.41$ years (blue) and $2.67$ years (red).
@@
As $\nu$ becomes smaller, the bifurcation diagrams becomes more and more complicated, with several nested limit cycle bubbles emerging from longer sequences of Hopf bifurcations. Figure 11 shows the stability switch curves for $\nu=1$, and the corresponding bifurcation diagram. In the left panel, we can see five different curves crossing the horizontal axis, giving rise to ten Hopf bifurcations. The last bifurcation, which corresponds to a stability switch from unstable to stable, is at approximately $\tau\approx 13.39$. The four rightmost bubbles of limit cycles are shown in the right panel. As before, the branch of limit cycles emerging from the smallest Hopf point is not shown, as it quickly diverges and the computation stops.
@@
Figure 13: Case $\nu=0$. Left: bifurcation diagram of $I$ (rightmost Hopf bifurcations) obtained with DDE-BIFTOOL, showing two regions of stability on different branches of periodic orbits. Right: stable limit cycles with different amplitude, obtained via long-term time integration using dde23.
@@ 3.1 The case with no boosting
$\dot{S}$ $=d(1-S)-\beta IS+\gamma e^{-d\tau}I(t-\tau)$
$\dot{I}$ $=\beta IS-(\gamma+d)I.$
@@
This corresponds to case 3 in the proof of Lemma 2. In our analysis with pertussis-like parameters, $J=(0.421,95.83)$. Since this interval is very large, we have covered only the region close to $\tau_{min}$ and $\tau_{max}$, where the $\omega^{\pm}$ are feasible and collide at both ends, see Figure 12 (top panels). There are $72$ possible stability switches, so in Figure 12 (bottom panels) we show only the first and last few $S_{n}^{\pm}(\tau)$ curves. The last switch is at the zero of $S_{35}^{-}$ at $\tau\approx 95.80$, Figure 12 (bottom right).
Thanks to the fact that the model has only one discrete delay, we performed the bifurcation analysis with the package DDE-BIFTOOL [12]. Figure 13 (left) shows the rightmost Hopf bifurcations and the two rightmost limit cycle bubbles. In this case, we did not detect bistability, but two stable limit cycles with different amplitude exist for parameter values that are in a relatively small interval around $\tau=95$ years. The right panel in Figure 13 shows the stable orbits in the $(S,I)$ plane, for $\tau=94$ (large amplitude) and $\tau=95.5$ (small amplitude). The convergence to the stable limit cycles was confirmed by performing a long-time integration of the system using the function dde23 for MATLAB.
Finally, we stress that, due to the large values of the delay at the stability switch when $\nu=0$, the numerical analysis with MatCont becomes challenging as large discretization indices are needed to obtain sufficient accuracy. Figure 14 compares the approximation error of the rightmost Hopf bifurcation detected by MatCont, for $\nu=3.2$ and $\nu=0$, depending on the degree $M$ of the collocation polynomial used in the approximation of the delay system in MatCont7p6. The right panel shows that a collocation degree of the order of $150$ is needed to approximate the Hopf bifurcation for $\nu=0$, which occurs at approximately $\tau\approx 95.80$. This is much larger than the collocation degree required to approximate the Hopf bifurcations for larger $\nu$ (left panel), which is of the order of $20$.
@@ 4 Discussion and conclusions
For parameter values describing a highly infectious disease like pertussis, our analysis uncovered a much more complex dynamical picture than previous ones, including catastrophic bifurcations and parameter regions with bistability of different attractors. While bistability in waning-boosting models has been observed previously (for instance [10] and [24]), our work is the first to analyse a delayed model and find evidence of bistability not only of the endemic equilibrium and a limit cycle, but also of two distinct stable limit cycles with different amplitude and period, for the same parameter values.
Catastrophic bifurcations and bistability have practical relevance in public health, as disturbances of the system’s state or parameters by non-pharmaceutical interventions can push the system to a different basin of attraction, hence potentially changing the long-term dynamics. Pertussis transmission, for instance, is highly influenced by waning-boosting dynamics, and post-pandemic outbreaks have sparked significant concern ([17]).
@@ Acknowledgments
The research of FS was supported by the Engineering and Physical Sciences Research Council via the Mathematical Sciences Small Grant UKRI170: “The dynamics of waning and boosting of immunity: new modelling and numerical tools”. FS is a member of the INdAM research group GNCS and of the UMI research group ‘Modellistica socio-epidemiologica’.

### v2

@@ 3 Numerical bifurcation analysis
To investigate the dynamical behaviour of the system beyond the equilibria, in this section we perform a numerical bifurcation analysis for the parameter set considered by [20] and [6], which is plausible with a highly infectious disease like pertussis. We take $\mathcal{R}_{0}=15$, $\gamma=17$ (year-1), corresponding to an average infectious period of 21 days, and per capita death rate $d=0.02$ (year-1), corresponding to an average lifetime of 50 years. The parameter $\beta$ is then fixed through (2) ($\beta=255.3$ year-1). We perform stability and bifurcation analyses varying the parameters $\tau$ and $\nu$, which represent the maximal duration of immunity in the absence of boosting and the boosting force, respectively. Since $\mathcal{R}_{0}>1$, the endemic equilibrium exists for all $\nu$ and $\tau$, with the susceptible population equal to $S^{*}=\mathcal{R}_{0}^{-1}=\frac{1}{15}$ at equilibrium.
@@
Figure 2: Left: stability of the endemic equilibrium of model (1) in the plane $(\nu,\tau)$. Solid lines represent Hopf bifurcation curves, with Hopf-Hopf (HH, $\bullet$) and Generalized Hopf (GH, $\circ$) bifurcations detected along the curves. The endemic equilibrium is unstable in the shaded area. Although not visible in the diagram due to the scaling of the axis, the lowest curves connect at a large value of $\nu$. The vertical dashed lines correspond to the parameters chosen for the one-parameter bifurcation diagrams in the following analyses. Note that an accurate approximation near $\nu=0$ requires large discretization indices and becomes computationally challenging, so we do not show it here. Right: close-up with the bifurcations numerically detected on different branches of limit cycles, delimiting their stability regions.
@@
Figure 2 shows the computed stability regions of the endemic equilibrium in the parameter plane $(\nu,\tau)$, which is consistent with the results presented by Barbarossa et al. [6, Figure 1]. The endemic equilibrium is stable outside the outermost boundary curve, and unstable inside (shaded region). Note that, as theoretically expected, the equilibrium is always stable on the axis $\tau=0$ (corresponding to an SIS-type model with individuals immediately susceptible upon recovery).
@@
GH points are often associated with the emergence of limit points of cycles (LPC) [19, Section 8.3.3] and hence with bistability of equilibria and periodic solutions. Similarly, HH points are associated with the emergence of Neimark–Sacker bifurcations (NS) and invariant tori [19, Section 8.6.3]. Figure 2 therefore suggests the emergence of rich and complex dynamics. We further explore such complex dynamics by studying the one-parameter continuation of the periodic orbits emerging from Hopf and their bifurcations, fixing $\nu=4.8$, 3.23.23.2, 222, 111, and varying the duration of immunity $\tau$ (along the dashed lines indicated in Figure 2).
Figure 3: Left: bifurcation diagram of $I$ varying $\tau$, for $\nu=4.8$, including the equilibrium curve with the detected Hopf bifurcations (H, $\bullet$) and the max/min values of the branch of limit cycles. Solid red lines indicate a stable equilibrium/limit cycle, dashed lines are unstable. Right: period of the limit cycle branch (years).
@@
Figure 3 shows the one-parameter bifurcation diagram with respect to $\tau$ for $\nu=4.8$ (left), and the corresponding estimated period along the branch of limit cycles (right). Four Hopf bifurcations are detected along the equilibrium curve, at approximately $\tau\approx 0.05$, 2.862.862.86, 3.153.153.15, and 4.594.594.59. The equilibrium is unstable between the first two points and between the last two points. A bubble of stable limit cycles exists for $\tau$ between 3.153.153.15 and 4.594.594.59 (both supercritical bifurcations), with a period between 2.2 and 2.8 years. When computing the limit cycle branch emerging from $\tau\approx 0.05$, the amplitude of the branch diverges quickly, and the numerical continuation stops. This branch is therefore not shown in the figure.
@@
As we can observe in Figure 5, the numerical computation of the rightmost eigenvalue crossings as $\tau$ varies (bottom left panel), aligns perfectly with the computation of possible stability switches, which are the zeros of $S^{\pm}_{n}(\tau)$, $n=0,1$ (top left panel). The zeros were calculated using (9) on the corresponding intervals $J^{+}$ and $J^{-}$. This also confirms Lemma 2, which implies that the endemic equilibrium will be stable for all $\tau>\tau_{max}$.
Figure 5: Top: graph of the stability switch functions $S_{n}^{\pm}(\tau)$ for $n=0,1,2$, where $\nu=4.8$ and 3.23.23.2, respectively. Bottom: corresponding real part of the rightmost eigenvalues along the equilibrium branch, varying $\tau$. Black dots indicate Hopf bifurcations.
To illustrate the emergence of bistability through the GH point, Figure 6 shows the one-parameter continuation for $\nu=3.2$. In this case, we still observe a bubble of limit cycles for $\tau$ between the Hopf points 2.112.112.11 and 5.375.375.37 (first supercritical, second subcritical). An LPC and an NS bifurcations are detected on the branch of limit cycles, and the limit cycles are stable between those bifurcations. Bistability occurs for values of $\tau$ between the Hopf and LPC bifurcations (i.e., for $\tau$ between 5.375.375.37 and 6.166.166.16), where a stable limit cycle coexists with the stable endemic equilibrium, so the asymptotic behaviour of the system depends on the initial condition.
@@
Figure 6: Left: bifurcation diagram of $I$ varying $\tau$, for $\nu=3.2$. Solid red lines indicate stable equilibrium/limit cycles; dashed lines indicate unstable. NS and LPC labels indicate Neimark-Sacker and limit point of cycles bifurcations, respectively. The blue shading shows the interval in which a stable equilibrium coexists with a stable limit cycle, for $\tau$ between 5.375.375.37 and 6.166.166.16. Right: period along the limit cycle branch (years).
Figure 7: Left: bistability of the positive equilibrium and the stable limit cycle for $\nu=3.2$ and $\tau=5.6$. Right: emergence of a stable torus via a Neimark–Sacker bifurcation, for $\nu=3.2$ and $\tau=4.2$.
@@
An even more complex dynamical picture is shown in Figure 8 for $\nu=2$, with two different closed limit cycle curves connecting different pairs of Hopf bifurcations, one with larger amplitude and larger period than the other. Each branch is stable in an interval of values between an LPC and an NS bifurcation. Two regions of bistability appear: one with a stable equilibrium coexisting with a stable limit cycle, and one with two different stable limit cycles. These regions are shaded in Figure 8. Figure 9 shows the orbits of the limit cycles in the small- and large-amplitude branches, as well as their time profiles. The figure highlights the differences in both amplitude and period. Following the LPC and NS curves for each bubble of limit cycles in the plane $(\nu,\tau)$ one can get some insights about regions of parameters where two stable periodic solutions coexist, or one stable periodic solution coexists with the stable endemic equilibrium. These curves are sketched in the right panel in Figure 2.
Figure 8: Left: bifurcation diagram of $I$ varying $\tau$, for $\nu=2$. Solid red lines indicate stable equilibrium/limit cycles; dashed indicate unstable. Right: period of the limit cycle along the branches. The figure shows two intervals of bistability: the blue shading indicates bistability of an equilibrium and limit cycle, while the red shading indicates bistability of two limit cycles differing in amplitude and period.
Figure 9: Orbits (left) and trajectories (right) of the stable coexisting limit cycles, for $\nu=2$ and $\tau=7$. Periods are approximately 1.411.411.41 years (blue) and 2.672.672.67 years (red).
@@
As $\nu$ becomes smaller, the bifurcation diagrams become more and more complicated, with several nested limit cycle bubbles emerging from longer sequences of Hopf bifurcations. Figure 11 shows the stability switch curves for $\nu=1$, and the corresponding bifurcation diagram. In the left panel, we can see five different curves crossing the horizontal axis, giving rise to ten Hopf bifurcations. The last bifurcation, which corresponds to a stability switch from unstable to stable, is at approximately $\tau\approx 13.39$. The four rightmost bubbles of limit cycles are shown in the right panel. As before, the branch of limit cycles emerging from the smallest Hopf point is not shown, as it quickly diverges and the computation stops.
@@
Figure 13: Case $\nu=0$. Left: bifurcation diagram of $I$ (rightmost Hopf bifurcations) obtained with DDE-BIFTOOL, showing two regions of stability on different branches of periodic orbits. Right: stable limit cycles with different amplitudes and periods, obtained via long-term time integration using dde23.
@@ 3.1 The case with no boosting
$\dot{S}(t)$ $=d(1-S(t))-\beta I(t)S(t)+\gamma e^{-d\tau}I(t-\tau)$
$\dot{I}(t)$ $=\beta I(t)S(t)-(\gamma+d)I(t).$
Similar models with a fixed immunity period and no boosting of immunity have been considered for instance by [15], [29], and more recently applied to COVID-19 transmission by [25].
@@
This corresponds to case 3 in the proof of Lemma 2. With our set of parameters, $J=(0.421,95.83)$. Since this interval is very large, we have covered only the region close to $\tau_{min}$ and $\tau_{max}$, where $\omega^{\pm}$ are feasible and collide at both ends, see Figure 12 (top panels). There are 727272 possible stability switches, so in Figure 12 (bottom panels) we show only the first and last few $S_{n}^{\pm}(\tau)$ curves. The last switch is at the zero of $S_{35}^{-}$ at $\tau\approx 95.80$, Figure 12 (bottom right).
[…]
Finally, we stress that, due to the large values of the delay at the stability switch when $\nu=0$, the numerical analysis with MatCont becomes challenging as large discretization indices are needed to obtain sufficient accuracy. Figure 14 compares the approximation error of the rightmost Hopf bifurcation detected by MatCont, for $\nu=3.2$ and $\nu=0$, depending on the degree $M$ of the collocation polynomial used in the approximation of the delay system in MatCont 7p6. The right panel shows that a collocation degree of the order of 150150150 is needed to approximate the Hopf bifurcation for $\nu=0$, which occurs at approximately $\tau\approx 95.80$. This is much larger than the collocation degree required to approximate the Hopf bifurcations for larger $\nu$ (left panel), which is of the order of 202020.
@@
Figure 15: Left: stability of the endemic equilibrium of model (1) in the plane $(\nu,\tau)$, with same parameters as in Figure 2, but the per capita mortality rate now fixed at $d=0.013$, corresponding to an expected lifespan of approximately 75 years. Right: stability of the endemic equilibrium in the plane $(d,\nu)$, for fixed $\tau=7$.
3.2 Impact of life expectancy
The previous analyses were carried out for fixed demographic parameters, and in particular for a rather short life expectancy (50 years). However, life expectancy is known to play an important role in the emergence of oscillatory dynamics: using a compartmental SIRWS model, [10] showed that increasing life expectancy can cause sustained oscillations in the presence of waning and boosting of immunity. With this motivation, we conclude this study with a brief discussion on the sensitivity of our stability results to the life expectancy of individuals.
The left panel of Figure 15 shows the stability and instability regions of the endemic equilibrium in the parameter plane $(\tau,\nu)$ (similarly as Figure 2) when the per capita death rate is $d=0.013$ years-1, corresponding to an average lifespan of approximately 757575 years. While the qualitative picture remains the same, decreasing the mortality rate $d$ enlarges the region of instability compared to Figure 2 (left). The destabilizing effect of increasing life expectancy is also evident from the right panel of Figure 15, which shows the stability and instability regions in the plane $(d,\nu)$, for fixed $\tau=7$ years and $\beta=255.3$ years-1, so that $\mathcal{R}_{0}$ varies with $d$ according to (2). We observe that, as $d$ decreases, the system undergoes a Hopf bifurcation with the emergence of stable periodic solutions. These results qualitatively agree with the analyses by [10] (see for instance Figure 4 in that reference).
@@ 4 Discussion and conclusions
For parameter values describing a highly infectious disease like pertussis, our analysis uncovered a much more complex dynamical picture than previous ones, including catastrophic bifurcations and parameter regions with bistability of different attractors. While bistability in waning-boosting models has been observed previously (for instance [10] and [24]), our work is the first to analyse a delayed model and find evidence of bistability not only of the endemic equilibrium and a limit cycle, but also of two distinct stable limit cycles differing in amplitude and period, for the same parameter values.
Catastrophic bifurcations and bistability have practical relevance in public health, as disturbances of the system’s state or parameters by non-pharmaceutical interventions can push the system to a different basin of attraction, hence potentially changing its long-term dynamics. Pertussis transmission, for instance, is highly influenced by waning-boosting dynamics, and post-pandemic outbreaks have sparked significant concern [17].
@@ Acknowledgments
The research of FS was supported by the Engineering and Physical Sciences Research Council via the Mathematical Sciences Small Grant UKRI170: “The dynamics of waning and boosting of immunity: new modelling and numerical tools”. FS is a member of the Computational Dynamics Laboratory (CDLab, University of Udine), of INdAM research group GNCS, of UMI research group Modellistica Socio-Epidemiologica, and of JUNIPER (Joint UNIversities Pandemic and Epidemiological Research).